A sample of radioactive material is initially found to have an activity of 115 decays/min. After 4 d 22 h, its activity is measured to be 63 decays/min.

1)

(a) Calculate the half-life of the material.
Answer #1

here ,

t = 4 + 22/24 = 4.917 days

as N = N0*(0.5)^(t/t0.5)

63 = 115*(0.5)^(4.917/t0.5)

(0.5)^(4.917/t0.5) = 0.548

taking ln

(4/t0.5) = 0.868

t0.5 = 5.664 days

the half life of the material is 5.664 days